OpenStudy (anonymous):

A local store owner pays her employees time-and-a-half for overtime. That means if an employee works more than 40 hours per week, the store will pay him 1.5 times his regular hourly wage of $6.50

OpenStudy (anonymous):

(a) Write and graph a piecewise function that illustrates an employee's weekly pay, W, as a function of the number of hours, h, that employee works

OpenStudy (anonymous):

remind me what a piecewise function is

OpenStudy (anonymous):

jk, found it http://www.mathsisfun.com/sets/functions-piecewise.html

OpenStudy (anonymous):

Maybe this will help you... someone tried this question before http://openstudy.com/study#/updates/516b671ce4b02ec89c5aac8b

OpenStudy (anonymous):

weekly pay, w will be our 'vertical axis' number of hours, h will be our 'horizontal axis' so the equation for the employee working 0-40 hours would be a straight line: y=mx+b or p=mh + b since regular wage is 6.5, then that is our slope, m we can figure out the 'b' term becuase it represens the y-intercept of the line. we know that if they work zero hours, they get paid zero dollars. so b=0 thus p = 6.5h + b that is the first part of the piece wise function

OpenStudy (anonymous):

I thought a piecewise function has < or > in it

OpenStudy (anonymous):

the second part will consist of the idea of being paid (1.5)6.5 dollars each hour over 40 hours. so this part of the function will start at 40 hours, and have a slope of 1.5*6.5 which is 6.5+3.25 = 9.75 = m so far the equation would be p=9.75h + b we now need to figure out what 'b' is. we know that at 41hrs we will be paid 9.75*41 = 399.75, from this knowledge we can figure out the point (41, 399.75), plug this into our equation: p=9.75h + b and solve for 'b'

OpenStudy (anonymous):

"I thought a piecewise function has < or > in it" u are right, we will get to this part next

OpenStudy (anonymous):

any questions so far on how to get the line equations for each function made so far?

OpenStudy (anonymous):

no, not yet

OpenStudy (anonymous):

so we have the two eqauations: f(h) = 6.5h for 0<= h <= 40 f(h) = 9.75h + b for 40<h *dont forget, 'b' needs to be solved for like i mentioned so here are our piecewise functions using < and >

OpenStudy (anonymous):

Ok I get that

OpenStudy (anonymous):

so how do we get b?

OpenStudy (anonymous):

take the equation: p=9.75h + b and plug in the point, (41, 399.75) then solve the equation for 'b'

OpenStudy (anonymous):

idk how to do that... p=9.75(41,399.75)+b? Also are you 100% on how to do this. I really need to get this right

OpenStudy (anonymous):

p=9.75(41,399.75)+b close, like this: 399.75 = 9.75(41) + b

OpenStudy (anonymous):

Oh I thought you meant like 41,399 thousand

OpenStudy (anonymous):

b=0?

OpenStudy (anonymous):

correct ^_^

OpenStudy (anonymous):

can you finish out the rest? or are there questions you have about what we did?

OpenStudy (anonymous):

Well just to clarify, what would I write as my answer for part a)?

OpenStudy (anonymous):

well, for the function we would have: f(h) = 6.5h for 0<= h <= 40 f(h) = 9.75h + b for 40<h as for the paragraph, you'll want to say something about how we made one function for 0 to 40 hours becuase the slope, or wage, was the same value. and we made a second function for h>40 hours becuase from here the wage changes from 6.5 to 9.75 becuase of the 'time and a half' for overtime.

OpenStudy (anonymous):

9.75h+b or 9.75h + 0? Also, are you 100% sure on these? Also again (sorry) I have to actually draw a graph

OpenStudy (anonymous):

oh, just do 9.75h

OpenStudy (anonymous):

yep, I am 100% sure, there is no other way possible we could do this

OpenStudy (anonymous):

how do I plot a graph? If you could show me or tell me how that would be amazing! I'll be back in an hour

OpenStudy (ranga):

"we know that at 41hrs we will be paid 9.75*41 = 399.75" may not be accurate. He will get paid the overtime rate only for the hours exceeding 40 hours a week. So if he works 41 hours, he will get paid: (6.50 * 40) + (1 * 9.75) = 260 + 9.75 = $269.75

OpenStudy (anonymous):

hey @ranga ! Are these piecewise functions correct? And how do I graph them?

OpenStudy (ranga):

W(h) = 6.5h for 0<= h <= 40 For hours exceeding 40 hours a week, the pay will be: (6.5 * 40) + (h - 40) * 9.75 = 260 + 9.75h - 390 = 9.75h - 130 W(h) = 9.75h - 130 for h > 40 The piecewise function is: W(h) = 6.5h for 0<= h <= 40 W(h) = 9.75h - 130 for h > 40 

OpenStudy (anonymous):

@ranga thanks for the clean up of my mess ^_^

OpenStudy (ranga):

Thanks to you @DemolisionWolf. You did most of the work.

OpenStudy (anonymous):

@ranga haha, I get sloppy the more time I sit on a problem!

OpenStudy (anonymous):

So using the piecewise functions as a base (W(h) = 6.5h for 0<= h <= 40 W(h) = 9.75h - 130 for h > 40 Determine how much an employee will be paid if he were to work 50 hours in a week.

OpenStudy (ranga):

h = 50 hours is more than 40 and so we have to use the second equation: W(h) = 9.75h - 130 for h > 40 Put h = 50 and calculate W(50)

OpenStudy (anonymous):

357.5

OpenStudy (ranga):

Yes, $357.50

OpenStudy (anonymous):

Oh ok, so that is my piecewise functions and graph? Thanks a lot both of you! I have my exam in 3 days so I'm trying to finish up the study guide. I have a few more questions today. Will you be available to help Ranga?

OpenStudy (ranga):

I will be here on and off as time permits. If you have any open questions when I am on I will try to answer.

OpenStudy (anonymous):

Got it! Oh and why do you have -130 when demo wolf had +b?

OpenStudy (ranga):

In an earlier reply I show how I derived the equation.